EduFarming
Your list · 0हिं

Type at least two letters. Press Esc to close.

↑↓ navigate↵ openTAB cycle filter

Conditional Probability & Bayes' Theorem · Question 59 of 125

Question 59

MediumOriginal question

AA and BB are independent with P(A)=0.3P(A) = 0.3 and P(B)=0.4P(B) = 0.4. What is P(A∪B)P(A \cup B)?

Options for question 59

Show the answer and explanation
CorrectNot quiteAnswer C

Correct answer: C

Explanation

  1. Independent: P(A∩B)=0.3×0.4=0.12P(A \cap B) = 0.3 \times 0.4 = 0.12.

  2. P(A∪B)=0.3+0.4−0.12=0.58P(A \cup B) = 0.3 + 0.4 - 0.12 = 0.58.

Worth noting

0.7 treats the events as mutually exclusive; independent events can happen together.

Concept
Probabilities with independent events
Typical time
90 s
Your time
—
The rule this testsIndependent events
Practise this chapter →

Ten questions from Conditional Probability & Bayes' Theorem, at your own pace.

Practise this chapter →

Ten questions from Conditional Probability & Bayes' Theorem, at your own pace.